OpenStudy (anonymous):

find the geometric mean of the pair 196 and 4

OpenStudy (anonymous):

geometric mean of two numbers a and b is given by \[b^2= ac\] where b-->geometric mean

OpenStudy (anonymous):

?

OpenStudy (anonymous):

therefore \[b = \sqrt{ac}\]

OpenStudy (anonymous):

thats the answer?

OpenStudy (anonymous):

i thought it was 28

OpenStudy (anonymous):

well,it is 28. You have to read first what I wrote before asking such questions

OpenStudy (anonymous):

@Sandiego19 , don't bother about the answer. Example sum. Find the geometric mean between 16 and 64 As I stated above . a= 16 , c=64 \[b =\sqrt{ac} = \sqrt{16*64} =4*8=32\] Now try your problem

OpenStudy (anonymous):

196 and 4 14x2 = 28

OpenStudy (anonymous):

yes @Sandiego19